# SimSUM – Simulated Benchmark with Structured and Unstructured Medical Records

### v3

@@ F.3 Neural-text
F.4 Neural-text-tab
For each evidence setting, we select the relevant set of tabular features and transform them into a numerical representation. We use a one-hot encoding for the categorical (binary or multiclass) features, and normalize the #days feature using the StandardScaler from sklearn. This tabular feature representation is then concatenated with the text representation we obtained in the previous baseline. Both are fed into the same architecture described in Section F.4, adapting the dimension of the input layer accordingly. For example, for the dyspnea classifier in the evidence setting all, the input dimension becomes $768+17$. We use the same hyperparameters as in the neural-text baseline to ensure a fair comparison. All other training details remain the same.

### v4

@@ E.3 Neural-text
E.4 Neural-text-tab
For each evidence setting, we select the relevant set of tabular features and transform them into a numerical representation. We use a one-hot encoding for the categorical (binary or multiclass) features, and normalize the days at home feature using the StandardScaler from sklearn. This tabular feature representation is then concatenated with the text representation we obtained in the previous baseline. Both are fed into the same architecture described in Section E.3, adapting the dimension of the input layer accordingly. For example, for the dyspnea classifier in the evidence setting all, the input dimension becomes $768+17$. We use the same hyperparameters as in the neural-text baseline to ensure a fair comparison. All other training details remain the same.